# More than Memes: A Multimodal Topic Modeling Approach to Conspiracy Theories on Telegram

### v2

@@ Results
The multimodal topics offers fewer aspects than the image setting, but the inclusion of textual information enhances interpretation and yields visually more diverse topics. We found one topic focused on actors, while others address the Israel-Gaza war, protests, and geopolitical issues, including references to historical events like the topic 1948 Nakba. Figure 4, row 1 shows images of Israel’s Prime Minister Benjamin Netanyahu. Accompanying text portrays him as a conspiratorial actor, often in antisemitic terms, illustrating the construction of enemy images and revealing connections strategies. One post labels him a “Zionist prostitute […] with the Zionist master lord Jacob Rothschild”, while another accuses Netanyahu and Zelensky of being “Jewish N…s” (omission in original) opposed to peace, and a third represents the strategy of providing evidence, referring to ‘internal sources’ claiming that Netanyahu supports using force against Iran’s nuclear program. Figure 4, row 2 shows examples of posts employing providing evidence and acts of disclosure: The first image, a black-and-white map purportedly from the radical Israeli group Irgun, is presented as evidence of a Israel’s alleged colonial agenda, with accompanying text claiming that it “is not an antisemitic invention that Israel is striving for a ‘Greater Israel”’. The second image features a long and cryptic QAnon-style text with phrases like “1st Map Palestine 1905” and “Balfour Declaration”. Notably, the post was assigned as outlier in the image setting, but assigned to the Israel Gaza topic in both the text and multimodal setting. Despite no visual similarity to the other two images in the row, apparently it’s textual dimension establishes a thematic link. This post exemplifies the authenticating strategy, presenting a screenshot of a post from the platform X by a conspiracy theorist with over 400K followers, an apparently influential figure. The third image is an Al Jazeera infographic, accompanied by text suggesting Gaza’s size could have tripled by 2020, possibly referencing then former US-American President Donald Trump’s alleged plan for a larger Palestinian state on the Sinai Peninsula. The use of an infographic conveys information in a factual manner, aimed at providing evidence.
@@
Cartoons in the context of the topic are employed for acts of disclosure, but also serve to create enemy images (Figure 4, row 3). The first cartoon shows a US-American puppet master with a Jewish ally, identified by national, religious, and political symbols. The text claims Hamas and IS were founded by the USA and Jews, who are further demonized through Nazi associations. The middle cartoon depicts Gaza as a prison under constant threat of bombs, while a Jewish soldier (marked by a Star of David on his helmet) portrays Israel (represented by its army), as victim. The cartoon claims to reveal Israel’s alleged evil and hypocrisy in falsely presenting itself as a victim. The accompanying text, “I love it”, reflects the poster’s support for the cartoon’s message. The cartoon on the right conveys a similar idea, highlighting double standards in how the war is perceived: it suggests that Palestinian civilian deaths are disregarded, while Israeli deaths receive more attention. The post’s text reinforces this, stating that “In the Anglo-Saxonian media it’s simple: in Israel there are many murders, and in Gaza there are many deaths”. This implies that ‘the media’ denies the deliberate killing of Palestinians while emphasizing Israeli casualties. The cartoon thus serves the strategies of discrediting mainstream accounts, accompanied by emotionalization.
@@ Conclusion
Our study applies multimodal topic modeling to German-language conspiracist Telegram data from October 2023, with additional analysis of a November 2024 dataset to evaluate temporal transferability. We demonstrate the utility of BERTopic for analyzing German-language text, image, and text-image Telegram posts, and integrate LLM-based topic annotation to increase efficiency and practical utility of the approach. Regarding predominant topics, we find that in October 2023, the Israel Gaza topic dominates across all topic models, while November 2024 focuses on German Politics, Ukraine War, and US Politics. Our analysis further reveals various image types, with photos and screenshots dominating, while memes play a minor role. The discovered visual types illustrate the diversity of visual content on Telegram, and also reflect varying text-image ratios: photos prioritize visuals, screenshots emphasize text, and memes represent an intermediary type. Our analysis of topic characteristics in different models reveals that topics are communicated differently: For instance, the largest part of the Israel Gaza topic group is present in text, while Protests rely heavily on images. Further, each model captures a substantial number of documents not captured by the other models, resulting in often small topic intersections. Topic symmetry is lower between text and image, and higher between image and multimodal models. Temporal transfer overall confirms our findings regarding the distribution of topic types, image types, symmetry and intersections of similar topics across models. Further, building on prior research on CT-related textual and visual meaning-making, we propose a novel framework for analyzing CT-related discursive strategies, comprising eight categories suited for the joint analysis of text and image data. Applying it to the Israel Gaza topic, we identify strategies such as providing evidence, authenticating, discrediting mainstream accounts, revealing connections, constructing enemy images, ingroup-outgroup framing, directives, and emotionalization, in text and image data.
@@ Discussion
Our study is subject to the following limitations: First, while we selected October 2023 to capture a period marked by heightened CT activity due to major crisis events (the Hamas attack on Israel and the Israel-Gaza war), this choice may have biased the dataset toward these events. To address this, we analyzed data from a later period, confirming findings such as screenshot predominance and low symmetry, even with thematic shifts. Second, our embedding model, pre-trained on descriptive image captions, is not optimized for noisy, user-generated texts with complex text-image relationships (e.g., commentative text-image pairs) (Theisen and Scheirer 2023, cf.). Fine-tuning CLIP on Telegram data and refining multimodal embedding strategies, such as adjusting text embedding weights, could improve topic quality. Third, the absence of standardized evaluation metrics for BERTopic models introduces subjectivity and limits comparability. Expanding the annotator pool in future studies could enhance validation reliability, and provide an even more solid gold standard for the evaluation of LLM-based performance.
@@
Future work could explore the reasons behind modality-specific differences in topic representation. While our study highlights notable variations in topic symmetry and intersections, a deeper investigation into the underlying causes is needed. Investigating the temporal dynamics of CT content across different periods, especially during less eventful times, could provide a broader understanding of how topics evolve over time. A systematic comparison of different platforms and content types could also offer valuable insights into the cross-platform dynamics of multimodal discourse. Insights into the relevance of screenshots open new avenues for studying sharing practices, information propagation and content dissemination, as well as the role of (manipulated) screenshots in disinformation on social media. Future work could further explore these practices, complementing existing research by including visual content. Finally, the proposed conceptual framework offers a foundation for both qualitative studies, such as multimodal discourse analysis, and quantitative studies, including annotation schema design for content analysis. It could be applied to diverse corpora, enhancing qualitative and quantitative analyses of how CTs are communicated across platforms through text, images, or their combination.
Ethics and Broader Perspective

### v3

@@ Results
The multimodal topics offers fewer aspects than the image setting, but the inclusion of textual information enhances interpretation and yields visually more diverse topics. We found one topic focused on actors, while others address the Israel Gaza war, protests, and geopolitical issues, including references to historical events like the topic 1948 Nakba. Figure 4 (row 1) shows images of Israel’s Prime Minister Benjamin Netanyahu. Accompanying texts portray him as a conspiratorial actor, often in antisemitic terms, illustrating the construction of enemy images and revealing connections strategies. One post labels him a “Zionist prostitute […] with the Zionist master lord Jacob Rothschild”, while another accuses Netanyahu and Zelensky of being “Jewish N…s” (omission in original) opposed to peace, and a third represents the strategy of providing evidence, referring to “internal sources”, claiming that Netanyahu supports using force against Iran’s nuclear program. Figure 4 (row 2) shows examples of posts employing providing evidence and acts of disclosure: the first image, a black-and-white map purportedly from the radical Israeli group Irgun, is presented as evidence of Israel’s alleged colonial agenda, with accompanying text claiming that it “is not an antisemitic invention that Israel is striving for a ‘Greater Israel”’. The second image features a long and cryptic QAnon-style text with phrases like “1st Map Palestine 1905” and “Balfour Declaration”. Notably, the post was assigned as outlier in the image setting, but assigned to the Israel Gaza topic in both the text and multimodal setting. Despite no visual similarity to the other two images in the row, apparently it’s textual dimension establishes a thematic link. This post exemplifies the authenticating strategy, presenting a screenshot of a post from the platform X by a conspiracy theorist with over 400K followers, an apparently influential figure. The third image is an Al Jazeera infographic, accompanied by text suggesting Gaza’s size could have tripled by 2020, possibly referencing then former US-American President Donald Trump’s alleged plan for a larger Palestinian state on the Sinai Peninsula. The use of an infographic conveys information in a factual manner, aimed at providing evidence.
@@
Cartoons in the context of the topic are employed for acts of disclosure, but also serve to create enemy images (Figure 4, row 3). The first cartoon shows a US-American puppet master with a Jewish ally, identified by national, religious, and political symbols. The text claims Hamas and IS were founded by the USA and Jews, who are further demonized through Nazi associations. The middle cartoon depicts Gaza as a prison under constant threat of bombs, while a Jewish soldier, marked by a Star of David on his helmet, portrays Israel, represented by its army, as victim. The cartoon claims to reveal Israel’s alleged evil and hypocrisy in falsely presenting itself as a victim. The accompanying text, “I love it”, reflects the poster’s support for the cartoon’s message. The cartoon on the right conveys a similar idea, highlighting double standards in how the war is perceived: it suggests that Palestinian civilian deaths are disregarded, while Israeli deaths receive more attention. The post’s text reinforces this, stating that “In the Anglo-Saxonian media it’s simple: in Israel there are many murders, and in Gaza there are many deaths”. This implies that ‘the media’ denies the deliberate killing of Palestinians while emphasizing Israeli casualties. The cartoon thus serves the strategies of discrediting mainstream accounts, accompanied by emotionalization.
@@ Conclusion
Our study applies multimodal topic modeling to German-language conspiracist Telegram data from October 2023, with additional analysis of a November 2024 data set to evaluate temporal transferability. We demonstrate the utility of BERTopic for analyzing German-language text, image, and text-image Telegram posts, and integrate LLM-based topic annotation to increase efficiency and practical utility of the approach. Regarding predominant topics, we find that in October 2023, the Israel Gaza topic dominates across all topic models, while November 2024 focuses on German Politics, Ukraine War, and U.S. Politics. Our analysis further reveals various image types, with photos and screenshots dominating, while memes play a minor role. The discovered visual types illustrate the diversity of visual content on Telegram, and also reflect varying text-image ratios: photos prioritize visuals, screenshots emphasize text, and memes represent an intermediary type. Our analysis of topic characteristics in different models reveals that topics are communicated differently: For instance, the largest part of the Israel Gaza topic group is present in text, while Protests rely heavily on images. Further, each model captures a substantial number of documents not captured by the other models, resulting in often small topic intersections. Topic symmetry is lower between text and image, and higher between image and multimodal models. Temporal transfer overall confirms our findings regarding the distribution of topic types, image types, symmetry and intersections of similar topics across models. Further, building on prior research on CT-related textual and visual meaning-making, we propose a novel framework for analyzing CT-related discursive strategies, comprising eight categories suited for the joint analysis of text and image data. Applying it to the Israel Gaza topic, we identify strategies such as providing evidence, authenticating, discrediting mainstream accounts, revealing connections, constructing enemy images, ingroup-outgroup framing, directives, and emotionalization, in text and image data.
@@ Discussion
Our study is subject to the following limitations: first, while we selected October 2023 to capture a period marked by heightened CT activity due to major crisis events such as the Hamas attack on Israel and the Israel-Gaza war, this choice may have biased the data set toward these events. To address this, we analyzed data from a later period, confirming findings such as screenshot predominance and low symmetry, even with thematic shifts. Second, our embedding model, pre-trained on descriptive image captions, is not optimized for noisy, user-generated texts with complex text-image relationships (e.g., commentative text-image pairs) (Theisen and Scheirer 2024a, cf.). Fine-tuning CLIP on Telegram data and refining multimodal embedding strategies, such as adjusting text embedding weights, could improve topic quality. Third, the absence of standardized evaluation metrics for BERTopic models introduces subjectivity and limits comparability. Expanding the annotator pool in future studies could enhance validation reliability, and provide an even more solid gold standard for the evaluation of LLM-based performance.
@@
Future work could explore the reasons behind modality-specific differences in topic representation. While our study highlights notable variations in topic symmetry and intersections, a deeper investigation into the underlying causes is needed. Investigating the temporal dynamics of CT content across different periods, especially during less eventful times, could provide a broader understanding of how topics evolve over time. A systematic comparison of different platforms and content types could also offer valuable insights into the cross-platform dynamics of multimodal discourse. Insights into the relevance of screenshots open new avenues for studying sharing practices, information propagation and content dissemination, as well as the role of (manipulated) screenshots in disinformation on social media. Future work could explore these practices further in order to complement existing research by including visual content. Our conceptual framework offers a foundation for both qualitative studies, such as multimodal discourse analysis, and quantitative studies, including annotation schema design for content analysis. It could be applied to diverse corpora, enhancing analyses of how CTs are communicated across platforms through text, images, or their combination.
Ethical Statement